OpenStudy (chosenmatt):

What is the first quartile (Q1) of the data set? 41, 33, 52, 67, 44, 38, 55, 62 A. 33 B. 39.5 C. 58.5 D. 67

OpenStudy (chosenmatt):

@iGreen

OpenStudy (igreen):

First we arrange from least to greatest: 33, 38, 41, 44, 52, 55, 62, 67 Now we split it in half and look at the first half: 33, 38, 41, 44 We take the two numbers in the middle, 38 and 41 and find the mean. 38 + 41 = ?

OpenStudy (chosenmatt):

79

OpenStudy (chosenmatt):

@iGreen

OpenStudy (igreen):

Yes, now divide by 2. 79 / 2 = ?

OpenStudy (chosenmatt):

what 2.79 and 2 come from?

OpenStudy (igreen):

We are finding the mean. And it's not 2.79, it's just 79. 79 came from adding 38 + 41 And finding the mean is adding the numbers up and dividing by how many numbers there are. There are 2 numbers, so we divide by 2. 79 / 2 = ?

OpenStudy (chosenmatt):

39.5

OpenStudy (igreen):

Yep, that's our answer.